Main INDEX
Monthly INDEX
PREV
NEXT
User name Seonho Choi/JPChen
Log entry time 00:01:47 on January17,2008
Entry number 226257
keyword=1/16 swing shift summary
1/16 Swing Shift Shift Summary (Seonho Choi, J. P. Chen, Yi Zhang)
------------------------------------------------------------------
16:00 waiting for Hall C to complete cryotarget cool-down..
17:20 Beam back to 50uA. Start production with solid targets
18:06 Hall C target cool-down completed, started to get power back to Hall A cryotarget
and start production on all targets.
21:23 change left P to 116.986 MeV, set by 21:28.
change right P to 175.561 MeV/c. settled by 22:40
22:35 change left P to 109.074 MeV/c, settled by 22:40
23:56 Data taking for CSR completed. Called to stop ebam.
Requested for controlled access for target warm-up.
24:00 turned spectrometer magnets to 0 current. Started cosmic runs.
Left Right p0_L p0_R Target Comments
---------------------------------------------------------------------------------------------------
5233 22810 0.14433 0.23812 Fe Production continues after ESR recovery
5234 22811 0.14433 0.23812 C12
5235 22812 0.14433 0.23812 DUMMY
5236 22813 0.13457 0.23812 DUMMY
5237 22814 0.13457 0.23812 C12
5238 22815 0.13457 0.23812 Fe
5239 22816 0.13457 0.23812 H2
5240 22817 0.13457 0.23812 Pb
5241 22818 0.13457 0.23812 He
5242 22819 0.14433 ------- He
5243 22820 0.14433 ------- Pb
5244 22821 0.14433 ------- H2
5245 22822 0.12547 ------- H2
5246 22823 0.12547 0.22202 Pb
5247 22824 0.12547 0.22202 He
5248 22825 0.12547 0.22202 LH2
5249 22826 0.12547 0.22202 DUMMY
5250 22827 0.12547 0.22202 C12
5251 22828 0.12547 0.22202 Fe
5252 22829 0.11699 ------- Fe
5253 22830 0.11699 ------- C12
5254 22831 0.11699 ------- DUMMY
5255 22832 0.11699 ------- LH2
5256 22833 0.11699 ------- Pb
5257 22834 0.11699 ------- He
5258 22835 0.10907 0.17556 He
5259 22836 0.10907 0.17556 Pb
5260 22837 0.10907 0.17556 LH2
5261 22838 0.10907 0.17556 DUMMY
5262 22839 0.10907 0.17556 C12
5263 22840 0.10907 0.17556 Fe